OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown.

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt[3]{x}^-6 }\]

OpenStudy (anonymous):

thats a 3 and a -6

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@wio

OpenStudy (debbieg):

So it's: \[\Large \dfrac{ 1 }{ \sqrt[3]{x}^{-6} }\] right?

OpenStudy (debbieg):

I would start by changing the root expression in the den'r to a rational number. Remember: the root is the den'r, the power is the num'r: \(\Large \sqrt[3]{x}=x^{1/3}\) and: \(\Large \sqrt[3]{x}^n=(x^{1/3})^n=x^{n/3}\)

OpenStudy (debbieg):

Use all that in the den'r, and then use the fact that: \(\Large \dfrac{1}{x^{-a}}=x^a\)

OpenStudy (anonymous):

-6 is inside the square root tho

OpenStudy (anonymous):

ohhhh so if \[\frac{ 1 }{ x^{-a} } = x{a} then the answer would be...\]

OpenStudy (anonymous):

then the answer would be.. hold on

OpenStudy (anonymous):

\[\sqrt[3]{x}^{6}\]

OpenStudy (debbieg):

It doesn't really matter if the power is inside the root or outside. :) Why? Because: \(\Large (x^{1/3})^n=(x^n)^{1/3}=x^{n\cdot (1/3)}=x^{n/3}\)

OpenStudy (anonymous):

ohhh i see i see

OpenStudy (debbieg):

You are correct in that you can take the den'r to the top (so it isn't a fraction anymore). However, what you have there is NOT rational exponent notation, so you aren't finished. You still have the root in radical notation. Change it to a rational exponent, and simplify. :)

OpenStudy (debbieg):

\(\Large \dfrac{ 1 }{ \sqrt[3]{x^{-6}} }=\sqrt[3]{x^{6}}\) Now put THAT into rational exponent form. \(\Large (x^{6})^?\)